7 customers enter a clothing store and are observed to determine if they will make a purchase. The store owner states that there is a 75% chance of a customer making a purchase. The random variable x represents a customer making a purchase. What is the probability that more than 3 customers make a purchase?
Answer #1

Solution

Given that ,

p = 75% = 0.75

1 - p = 0.25

n = 7

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X > 3) = P(X = 4) +  P(X = 5) + P(X = 6) + P(X = 7)

= (7 C 4) * 0.754 * (0.25)3 + (7 C 5) * 0.755 * (0.25)2 + (7 C 6) * 0.756 * (0.25)1 + (7 C 7) * 0.757 * (0.25)0

= 0.9294

Probability = 0.9294